OpenStudy (anonymous):

I have this math equation that I need help with, if anyone can help me. (25x^3)^2/3 the 2/3 is in a fraction form ----------- (125xy^2)^1/3 the 1/3 is in a fraction form too..

OpenStudy (anonymous):

So in total does it look like \[\frac{(25x^3)^{\frac{2}{3}}}{(125xy^2)^{\frac{1}{3}}}\]?

OpenStudy (anonymous):

Or are the two parts equal to each other?

OpenStudy (anonymous):

that looks exactly..

OpenStudy (anonymous):

the same

OpenStudy (anonymous):

Alright. So is the objective to simplify the fraction?

OpenStudy (anonymous):

yeah, i don't know where to start...

OpenStudy (angela210793):

my internet doesn't want me to help u....I'm trying to upload the pic but nothing :(:(:(

OpenStudy (anonymous):

\[(\sqrt{25})^{3}(x ^{3})^{2/3}\] Numerator will be 125x^2

OpenStudy (anonymous):

It's definitely a tough one. Let me see if I can make it look a bit nicer \[\frac{(25x^3)^{\frac{2}{3}}}{(125xy^2)^{\frac{1}{3}}}=\left[\frac{(25x^3)^2}{125xy^2}\right]^{\frac{1}{3}}\]

OpenStudy (angela210793):

i expressed 25 as 5^2 and 125 as 5^3 and then subtracted the exponents

OpenStudy (anonymous):

\[(\sqrt[3]{125})*(\sqrt[3]{x})*(\sqrt[3]{y ^{2}})\]

OpenStudy (anonymous):

I don't think you can subtract the exponents when they're inside brackets taken to different powers. You'd have to apply the square to the numerator before you could do that.

OpenStudy (angela210793):

i have expressed them the same with base=5 and if a^x/a^y=a^(x-y) isn't it?

OpenStudy (anonymous):

Yes, but that is not quite the case here. You have \[\frac{(a^x)^{\frac{2}{3}}}{(a^y)^{\frac{1}{3}}}\] You need to bring that 2 inside, so it's \[\frac{(a^{2x})^{\frac{1}{3}}}{(a^y)^{\frac{1}{3}}}\]

OpenStudy (angela210793):

yea :):) I'm not gonna write anymore :) u keep going with ur solution cause it is right as well :):):):) sorry :)

OpenStudy (anonymous):

Anyway, if you're still around monika, how did you get on with that?

OpenStudy (anonymous):

I have no idea, what to do...

OpenStudy (anonymous):

Dalvoron, can u show me again

OpenStudy (anonymous):

Alright, well let's start with the last thing I posted \[\left[\frac{(25x^3)^2}{125xy^2}\right]^{\frac{1}{3}}\] The first thing I want you to do is apply that power of 2. Just look at the \((25x^3)^2\) for now, and try to apply that square.

OpenStudy (anonymous):

Use the rule \((a b)^2 = (a^2b^2)\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

What do you get when you do that?

OpenStudy (anonymous):

ok, hang on.. still working on it..

OpenStudy (anonymous):

do i work with 25x^3 becomes as what

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

but what happens with ^2

OpenStudy (anonymous):

You have to apply it to the 25, and the \(x^3\)

OpenStudy (anonymous):

ok, so its (25^2) (x^3)^2

OpenStudy (anonymous):

Yep, and what does that equal?

OpenStudy (anonymous):

625 but i don't know how to solve (x^3)^2

OpenStudy (anonymous):

The 625 is right. Use the rule that \(x^a)^b = x^{(a)(b)}\)

OpenStudy (anonymous):

ok, so its.. x^6

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Alright, so now we have \[\left[\frac{(25x^3)^2}{125xy^2}\right]^{\frac{1}{3}}=\left[\frac{625x^6}{125xy^2}\right]^\frac{1}{3}\]

OpenStudy (anonymous):

Since everything is inside one pair of square brackets, the next step is to ignore them for the moment, and focus on the fraction \[\frac{625x^6}{125xy^2}\] How would you simplify that?

OpenStudy (anonymous):

do i divide the 625 by 125

OpenStudy (anonymous):

Yep, that's one of the two things you do at this stage.

OpenStudy (anonymous):

What does the fraction look like after you do that division?

OpenStudy (anonymous):

i got 5x^6 ----- xy^2

OpenStudy (anonymous):

That's right! What else could you do to simplify the fraction?

OpenStudy (anonymous):

no clue

OpenStudy (anonymous):

Let's say you had just the xs there, so your fraction was \[\frac{x^6}{x}\] Could you simplify that?

OpenStudy (anonymous):

can it be 2, 6 divided by 2 and the answer is 3

OpenStudy (anonymous):

Not that, no. I'll try to write it differently, maybe it will help. \[\frac{x^6}{x} = \frac{x \times x \times x \times x \times x \times x}{x}\]

OpenStudy (anonymous):

it becomes x^5 cause one set of the x's cancels out..

OpenStudy (anonymous):

Exactly So now your fraction is \[\left[\frac{5x^5}{y^2}\right]^\frac{1}{3}\]

OpenStudy (anonymous):

As far as I can tell, it doesn't get any simpler than that.

OpenStudy (anonymous):

this equation looks simple for you but not for me... so, from here what else do I do...

OpenStudy (anonymous):

What I mean is that I think that is the end of the question.

OpenStudy (anonymous):

lol.. oh, i knew that..

OpenStudy (phi):

Is this multiple choice?

OpenStudy (anonymous):

what..

OpenStudy (phi):

The answer Dal helped you with is good. But sometimes, with multiple choice answers, it won't be there... there are other ways to write this expression. For example \[(5 x^{5}y^{-2})^{\frac{1}{3}}\]